# Mind the Gap: a Spectral Analysis of Rank Collapse and Signal Propagation in Transformers

### v1

@@ A.1 Proofs
Proof of Proposition 1.
Let us show that the attention matrix $\mathbf{A}_{1}(\mathbf{X}_{0})$ satisfies Definition 1.1 by demonstrating that the random variables
@@
Since the key and query matrices are initialised as Gaussian i.i.d. matrices and the input data $\mathbf{X}_{0}$ is isometric, $\widetilde{\mathbf{W}}^{Q}\coloneqq\mathbf{X}_{0}\mathbf{W}_{1}^{Q}$ and $\widetilde{\mathbf{W}}^{K}\coloneqq\mathbf{X}_{0}\mathbf{W}_{1}^{K}$ follow the same distribution as $\mathbf{W}_{1}^{Q}$ and $\mathbf{W}_{1}^{K}$. Each $Z_{i,j}$ can be written as the exponential of the inner product between the $i$-th row of $\widetilde{\mathbf{W}}^{Q}$ and the $j$-th row of $\widetilde{\mathbf{W}}^{K}$, thus $Z_{i,j}$ are i.i.d. and we only need to prove that $\mathbb{E}|Z_{1,1}|^{4}<\infty$. Let us define
@@
as long as $\frac{4}{\sqrt{d_{qk}}}<1$, i.e. $d_{qk}>16$. ∎
@@ A.2 Lemmas
holds with probability at least $1-2\exp\big{[}-C^{\prime}d^{q}\min(t^{2},t^{2/q})\big{]}$, i.e. with overwhelming probability [Tao12, Definition 1.1.2]
@@
Consider $p$ Random Markov matrices $\mathbf{A}_{1},\dots,\mathbf{A}_{p}\in\mathbb{R}^{T\times T}$ as defined in 1.1, and let $\mathbf{1}_{T\times T}$ be the matrix full of ones. Then, almost surely,
@@
which yields, almost surely, $\lim s_{1}(\mathbf{A}_{p}\dots\mathbf{A}_{1})\leq 1$. The converse inequality is an immediate consequence of the closure of the set of Random Markov matrices under matrix multiplication, which gives $\lambda_{1}(\mathbf{A}_{p}\dots\mathbf{A}_{1})=1$, combined with $s_{1}(\mathbf{A}_{p}\dots\mathbf{A}_{1})\geq|\lambda_{1}(\mathbf{A}_{p}\dots\mathbf{A}_{1})|$. Hence, almost surely, $\lim s_{1}(\mathbf{A}_{p}\dots\mathbf{A}_{1})=1$.
@@
Let $\mathbf{A}_{1},\dots,\mathbf{A}_{p}\in\mathbb{R}^{T\times T}$ be independent Random Markov matrices as defined in 1.1 and $\mathbf{W}_{1}\in\mathbb{R}^{T\times d}$, $\mathbf{W}_{2},\dots,\mathbf{W}_{q}\in\mathbb{R}^{d\times d}$ be independent Gaussian matrices with i.i.d. $\mathcal{N}(0,1)$ entries. Then, for large enough $T$ and $d$ with fixed $\gamma=T/d\in(0,1]$, the event
@@
Let $\mathbf{A}_{1},\dots,\mathbf{A}_{p}\in\mathbb{R}^{T\times T}$ be Random Markov matrices as defined in 1.1 and $\mathbf{W}_{1}\in\mathbb{R}^{T\times d}$, $\mathbf{W}_{2}\dots,\mathbf{W}_{q}\in\mathbb{R}^{d\times d}$ be independent Gaussian matrices with i.i.d. $\mathcal{N}(0,1)$ entries. Then, for $T$ and $d$ large enough,
@@
Let $\mathbf{A}\in\mathbb{R}^{T\times T}$ be a Random Markov matrix, and let $\mathbf{A}^{\perp}\coloneqq\mathbf{A}-T^{-1}\mathbf{1}_{T\times T}$. Then, almost surely, the empirical singular value distribution of $T^{1/2}\mathbf{A}^{\perp}$ weakly converges to the quartercircular law as $T\to\infty$, i.e.
@@
The distribution in question $\mathcal{M}$ is the limiting squared singular value distribution of a product of rectangular independent Gaussian matrices, whose general moments are worked out in [AIK13, equation 58]. Simple algebraic manipulations lead to our result. ∎
A.3 Supplementary data
A.3.1 Implementation details
@@
The default model consists of a stack of single-head attention layers, with an optional LayerNorm inserted between them (denoted by “+ LN” in the legend) after receiving an optional skip connection from the previous layer (denoted by “+ skip” in the legend). When both options are enabled simultaneously, the configuration is referred to as “+ skip + LN”. By single-head, we mean that only one attention mechanism is computed, applied to the values and then multiplied by a matrix $\mathbf{W}_{h}$ which is initialised as the identity matrix and optimised during training.
@@
Training.
Given some isometric $\mathbf{X}_{0}$ input data, the goal is to learn the entrywise Heaviside function, a non trivial task due to the function’s nonlinearity. To achieve this, we train a series of attention-only transformer encoders on a mean squared error (MSE) loss, optimised with Adam. We conduct an extensive grid search over the learning rate $\lambda\in\{1,3,5\}\times{10}^{-\{1,2,3,4,5\}}$. Each experiment is run $5$ times, and the learning rate that results in the best average training performance for each configuration, as shown in the plots, is selected. The figures display the training loss with respect to training steps, i.e. the number of gradient descent updates. A “no training” label is shown when no training progress is made after $50$ training epochs, despite tuning the learning rate. We train on a set of $50$ data points, each of size $T\times d$, with $T=d=500$ to ensure we are in the large width regime that our theoretical framework presupposes.
A.3.2 Additional experiments
(a) $L=2$
(b) $L=4$
(c) $L=6$
(d) $L=3$
(e) $L=5$
(f) $L=7$
Figure 9: In a transformer with key-query attention, the gradient norm explodes in width at a rate that worsens with increasing depth $L$, exceeding the growth of $T^{L-1}$ predicted by our analysis.

### v2

@@ Appendix A Appendix
The Appendix is organised as follows:
  • Section A.1 presents the proofs of the paper’s propositions.
  • Section A.2 provides auxiliary lemmas along with their proofs that support the previous section.
  • Section A.3 explores dynamical isometry in transformers, questioning their ability to propagate information faithfully through attention layers once the problematic outliers have been removed.
  • Section A.4 outlines comprehensive experimental details, ensuring reproducibility.
@@ A.1 Proofs
Proof of Theorem 1.
Let us show that the attention matrix $\mathbf{A}(\mathbf{X}_{0})$ satisfies Definition 3.1 by demonstrating that the random variables
@@
Since the key and query matrices are initialised as Gaussian i.i.d. matrices and the input data $\mathbf{X}_{0}$ is isotropic, $\widetilde{\mathbf{W}}^{Q}\coloneqq\mathbf{X}_{0}\mathbf{W}_{1}^{Q}$ and $\widetilde{\mathbf{W}}^{K}\coloneqq\mathbf{X}_{0}\mathbf{W}_{1}^{K}$ follow the same distribution as $\mathbf{W}_{1}^{Q}$ and $\mathbf{W}_{1}^{K}$. Each $Z_{i,j}$ can be written as the exponential of the inner product between the $i$-th row of $\widetilde{\mathbf{W}}^{Q}$ and the $j$-th row of $\widetilde{\mathbf{W}}^{K}$, thus $Z_{i,j}$ are i.i.d. and we only need to prove that $\mathbb{E}|Z_{1,1}|^{4}<\infty$. Let us define
@@
as long as $\frac{4}{\sqrt{d_{qk}}}<1$, i.e. $d_{qk}>16$.
It is now time to invoke an important result, that we first state.
Theorem 8 (Spectral gap in i.i.d. Markov matrices, [BCC11]).
Let $\mathbf{A}\in\mathbb{R}^{T\times T}$ be an i.i.d. Markov matrix whose underlying distribution has variance $\sigma_{A}^{2}$. Then, $\lambda_{1}(\mathbf{A})=1$ and almost surely,
[…]
Having shown that $\mathbf{A}(\mathbf{X}_{0})$ forms an i.i.d. Markov matrix since the $Z_{i,j}$’s are i.i.d. with finite fourth moment, the spectral gap as described in Theorem 1 follows immediately. ∎
@@ A.2 Lemmas
holds with probability at least $1-2\exp\big{[}-C^{\prime}d^{q}\min(t^{2},t^{2/q})\big{]}$, i.e. with overwhelming probability [Tao12]. ∎
@@
Consider $p$ i.i.d. Markov matrices $\mathbf{A}_{1},\dots,\mathbf{A}_{p}\in\mathbb{R}^{T\times T}$ as defined in 3.1, and let $\mathbf{1}_{T\times T}$ be the matrix full of ones. Then, almost surely,
@@
which yields, almost surely, $\lim s_{1}(\mathbf{A}_{p}\dots\mathbf{A}_{1})\leq 1$. The converse inequality is an immediate consequence of the closure of the set of i.i.d. Markov matrices under matrix multiplication, which gives $\lambda_{1}(\mathbf{A}_{p}\dots\mathbf{A}_{1})=1$, combined with $s_{1}(\mathbf{A}_{p}\dots\mathbf{A}_{1})\geq|\lambda_{1}(\mathbf{A}_{p}\dots\mathbf{A}_{1})|$. Hence, almost surely, $\lim s_{1}(\mathbf{A}_{p}\dots\mathbf{A}_{1})=1$.
@@
Let $\mathbf{A}_{1},\dots,\mathbf{A}_{p}\in\mathbb{R}^{T\times T}$ be independent i.i.d. Markov matrices as defined in 3.1 and $\mathbf{W}_{1}\in\mathbb{R}^{T\times d}$, $\mathbf{W}_{2},\dots,\mathbf{W}_{q}\in\mathbb{R}^{d\times d}$ be independent Gaussian matrices with i.i.d. $\mathcal{N}(0,1)$ entries. Then, for large enough $T$ and $d$ with fixed $\gamma=T/d\in(0,1]$, the event
@@
Let $\mathbf{A}_{1},\dots,\mathbf{A}_{p}\in\mathbb{R}^{T\times T}$ be i.i.d. Markov matrices as defined in 3.1 and $\mathbf{W}_{1}\in\mathbb{R}^{T\times d}$, $\mathbf{W}_{2}\dots,\mathbf{W}_{q}\in\mathbb{R}^{d\times d}$ be independent Gaussian matrices with i.i.d. $\mathcal{N}(0,1)$ entries. Then, for $T$ and $d$ large enough,
@@
Let $\mathbf{A}\in\mathbb{R}^{T\times T}$ be an i.i.d. Markov matrix, and let $\mathbf{A}^{\perp}\coloneqq\mathbf{A}-T^{-1}\mathbf{1}_{T\times T}$. Then, almost surely, the empirical singular value distribution of $T^{1/2}\mathbf{A}^{\perp}$ weakly converges to the quartercircular law as $T\to\infty$, i.e.
@@
The distribution in question $\mathcal{M}$ is the limiting squared singular value distribution of a product of rectangular independent Gaussian matrices, whose general moments are worked out in [AIK13]. Simple algebraic manipulations lead to our result. ∎
A.3 Can transformers achieve dynamical isometry?
In Section 3, we have established (i) the existence of an outlier eigenvalue/singular value in the spectrum of softmax-based attention matrices, and (ii) that removing this outlier helps with rank collapse and exploding gradients. In the absence of the outlier, we can take a further step to analyse the bulk of the spectra of the network’s token-wise covariance and input-output Jacobian—two quantities that have been shown to play a key role in signal propagation as we will see below. To this end, we will make use of free probability theory.
Free probability.
Let $\mathbf{X}^{\perp}_{\ell}=\mathbf{A}^{\perp}_{\ell}\mathbf{X}^{\perp}_{\ell-1}\mathbf{W}^{V}_{\ell}$ be the signal at layer $\ell$ in our modified model (16) and $\mathbf{\Sigma}^{\perp}_{\ell}=\mathbf{X}^{\perp}_{\ell}{\mathbf{X}^{\perp}_{\ell}}^{\top}\in\mathbb{R}^{T\times T}$ be its covariance matrix. Let the underlying i.i.d. Markov matrices $\mathbf{A}_{\ell}$ have variance $\sigma_{A}^{2}$ and $\mathbf{W}^{V}_{\ell}$ have i.i.d. $\mathcal{N}(0,\sigma_{V}^{2})$ entries. Let $\boldsymbol{\Sigma}^{\perp}_{0}=\mathbf{I}$ and $\mathcal{D}_{\ell}$ be the limiting singular value distribution of $\boldsymbol{\Sigma}^{\perp}_{\ell}$. Then the mean and variance of $Z\sim\mathcal{D}_{\ell}$ are given by
$\mathbb{E}(Z)$ $=(\sigma_{A}\sigma_{V}/\sqrt{\gamma})^{2\ell},$ (35)
where $\gamma\coloneqq\frac{T}{d}\in(0,1]$.
It is evident from the above proposition that simultaneously controlling both the mean and variance of $\mathcal{D}_{\ell}$ is not feasible. Model (16) does not have enough hyperparameters to achieve this balance. Indeed, to prevent the mean from growing or shrinking exponentially with depth, the product $\sigma_{A}\sigma_{V}$ must equal $\sqrt{\gamma}$. However, this constraint leads to the variance increasing linearly with $\ell$.
The goal is to ensure that the spectral energy of the Jacobian concentrates around 1, thereby minimising distortion of the input space geometry—a property often referred to as the dynamical isometry in the literature (see [PSG17]). For our model (16), it is straightforward to show
where $\otimes$ denotes the Kronecker product. Proposition  12 describes the moments of the limiting squared singular value distribution of $\mathbf{J}_{\ell}$.
Let $\mathbf{X}^{\perp}_{\ell}=\mathbf{A}^{\perp}_{\ell}\mathbf{X}^{\perp}_{\ell-1}\mathbf{W}^{V}_{\ell}$ be the signal at layer $\ell$ in our modified model (16). Let the underlying i.i.d. Markov matrices $\mathbf{A}_{\ell}$ have variance $\sigma_{A}^{2}$ and $\mathbf{W}^{V}_{\ell}$ have i.i.d. $\mathcal{N}(0,\sigma_{V}^{2})$ entries. Let $\mathcal{D}_{\ell}$ be the limiting distribution of the squared singular values of $\mathbf{J}_{\ell}\coloneqq\partial\mathbf{X}^{\perp}_{\ell}/\partial\mathbf{X}_{0}$ [^5] In a minor abuse of notation, we may write $\partial\mathbf{X}^{\perp}_{\ell}/\partial\mathbf{X}_{0}$ as a shorthand for $\partial\,\mathrm{vec}(\mathbf{X^{\perp}_{\ell}})/\partial\,\mathrm{vec}(\mathbf{X}_{0})$.. Then the mean and variance of $Z\sim\mathcal{D}_{\ell}$ are given by
$\mathbb{E}(Z)$ $=(\sigma_{A}\sigma_{V})^{2\ell},$ (40)
$\mathrm{Var}(Z)$ $=\ell(\ell+2)(\sigma_{A}\sigma_{V})^{4\ell}.$ (41)
A.4 Implementation details
@@
The default model consists of a stack of single-head attention layers, with an optional LayerNorm inserted between them (denoted by “+ LN” in the legend) after receiving an optional skip connection from the previous layer (denoted by “+ skip” in the legend). When both options are enabled simultaneously, the configuration is referred to as “+ skip + LN”. By single-head, we mean that only one attention mechanism is computed, applied to the values and then multiplied by a matrix $\mathbf{W}_{h}$ which is initialised as the identity matrix.
@@
$\mathbf{A}^{\text{no outliers}}=\mathbf{A}-\sum_{i=1}^{r}s_{i}u_{i}v_{i}^{\top},$
where $\mathbf{A}=\mathbf{U}\mathrm{diag}\big{(}s_{1}(\mathbf{A}),\dots,s_{n}(\mathbf{A})\big{)}\mathbf{V}^{\top}$. The cut off threshold $r$ is chosen such that it maximises the difference between consecutive singular values, i.e.
$r\coloneqq\operatorname*{arg\,max}_{1\leq i\leq n-1}\,(s_{i}-s_{i+1}).$